OpenStudy (angel_kitty12):

Find the inverse of the function f(x)=7/x-3

OpenStudy (angel_kitty12):

\[f(x)=\frac{ 7 }{ x }-3\]

OpenStudy (angel_kitty12):

I got halfway before getting confused

OpenStudy (angel_kitty12):

swap x for y f(x)=7/y-3 add three to both sides 7/y=x+3

OpenStudy (angel_kitty12):

then i didnt knwo how to solve the fraction from there

OpenStudy (lowkey.s):

okay so to find a inverse function, interchange the variebles and solve for y.

OpenStudy (angel_kitty12):

I did that but I just got stuck on how to solve the last step. I got until \[\frac{ 7 }{ y }=x+3\]

OpenStudy (lowkey.s):

okay so